# Climate Physical Risk Assessment in Asset Management

### v1

@@ Abstract
Climate-related phenomena are increasingly affecting regions worldwide, manifesting as floods, water scarcity, and heat waves, which can significantly impair companies’ assets and productivity. We develop a framework, based on the Vasicek model for credit risk, that introduces downward jumps due to climate phenomena in a company asset’s dynamics. These negative shocks are designed to mirror the negative effect of extreme climate events. We explore various possibilities for the distribution of jumps and provide a comprehensive characterization of the model and its calibration process that relies on companies’ asset intensity and geographical exposure. We utilize this new multivariate firm value model framework with jumps to understand the influence of climate related extreme events on different portfolios’ expected and unexpected losses. The results suggest introducing additional safe capital to offset the losses resulting from physical climate risks, particularly for high asset intensity sectors.
()($\ddagger$)( ‡ ) Politecnico di Milano, Department of Mathematics, 32 p.zza L. da Vinci, Milano
()(*)( ∗ ) Arca Fondi SGR, 3, Via Disciplini, Milano
@@
Climate Physical Risk Assessment in Asset Management
@@ 1 Introduction
Physical risk emerging from climate change will play a key role in the asset management and banking industry. It comprises both acute risks stemming from climate-related hazards (e.g., from extreme weather events) and chronic risks (e.g., from sea-level rise) from long-run climate impacts (Pankratz et al. 2023, Bressan et al. 2024). In response, central banks and supervisory authorities globally expect financial institutions with material climate-related and environmental exposures to integrate physical climate risks into their risk management frameworks. [^1] See e.g., the ECB https://www.ecb.europa.eu/press/economic-bulletin/focus/2023/html/ecb.ebbox202302_06~0e721fa2e8.en.html, the FED https://www.federalreserve.gov/publications/files/csa-exercise-summary-20240509.pdf, and the Basel Committee on Banking Supervision perspectives https://www.bis.org/bcbs/publ/d560.pdf. We propose a new and simple model to include this risk in asset management decisions, using a multivariate approach. We model physical risk impact on stock portfolios, aggregating the firm-level effects of such risk. To the best of our knowledge, this is the first model in the literature that exhibits these characteristics and that can be flexibly applied to any stocks portfolio using widely available data.
@@
Physical risk affects directly firms’ tangible assets (e.g. production plants or equipment) as well as operating costs, such as relocating and insurance costs. These shocks result in lower profits and reduced repayment capacity, which are then transmitted to the quoted stock prices (Ginglinger and Moreau 2023). To mirror this cascading effect we base our approach on Vasicek (1987), a multivariate generalization of the simple Merton (1974) Firm Value Model (FVM) that considers a single market risk factor. In this framework, the firms’ asset value is modeled with a geometric Brownian motion, while equity is modeled as a European call option on the firm’s asset value, with the company’s debt as strike price. We propose a modification of the Vasicek (1987) model that incorporates jumps to capture the impact of climate events which adversely affect a firm’s asset value. The jump process frequency and size depend on the firm’s characteristics such as its geographic exposure and how much of its revenues are linked to physical assets.
@@
In our study, we follow an efficient and versatile approach for the calibration. It is based on widely available data that covers most quoted stocks worldwide and enables low-cost implementation by financial institutions. Specifically, we focus on the exposure of physical assets to climate risks. We utilize the publicly available ND-GAIN (Notre Dame Global Adaptation Initiative) Country Index to assess risk based on geographical location and incorporate balance sheet data to estimate the proportion of a company’s assets exposed to such risks. [^2] See https://gain.nd.edu/assets/522870/nd_gain_countryindextechreport_2023_01.pdf. The ND GAIN index is a widely used benchmark for climate risk in the literature (see e.g., Russo et al. 2019, Dechezleprêtre et al. 2022). However, our model is very flexible since it can be calibrated starting from different definitions of climate losses such as the ones computed by Bressan et al. (2024) for Mexican companies or from private data providers such as those listed in Hain et al. (2022) or the CRIS methodology discussed in Ginglinger and Moreau (2023).
@@
Once the extended Vasicek (1987) model is calibrated, we can utilize it to compute different climate-related risk metrics for a portfolio. In our analysis, we focus on computing an add-on for the Value-at-Risk (VaR) of a portfolio that accounts for a climate-stressed scenario [^3] Our methodology can be easily extended to compute an add-on for the Expected Shortfall (ES). We evaluate the evolution of such metrics for different time horizons. This is particularly relevant from the perspective of financial institutions and supervisors, as it enables the incorporation of physical risks into their risk management practices building upon their existing frameworks.
@@
The majority of the literature on climate risk impact on portfolios focuses on transition risk and not on physical risk. Transition risk stems from the transition to a low-carbon economy. It includes risks created by mitigation and adaptation policies, emerging clean technologies, and behavioral changes of consumers and investors (Bolton and Kacperczyk 2023). Le Guenedal and Tankov (2024) utilize a FVM that accounts for the uncertainty of the carbon tax trajectory while Kölbel et al. (2024) show that, in a FVM approach, CDS spread are affected by transition and physical risk news. De Angelis et al. (2023) discuss the influence of green investors on long-run emissions trajectories, finding that tighter climate regulations and climate-related technological innovations abate carbon emissions, in contrast to uncertainty regarding future climate. Bolton and Kacperczyk (2023) point out that transition risk is not significantly related to different exposures to physical risk.
@@
We propose an alternative approach that strikes a balance between the two extrema discussed above: Dietz et al. (2016) estimate only a global loss, while Bressan et al. (2024) provide highly detailed asset-level estimations for a single country, which are challenging to apply for portfolio risk assessment. Our model, in contrast, can be easily applied to any portfolio of assets for which climate loss estimates and balance sheet data are available, providing various climate-related risk metrics at both the portfolio and asset levels. We also outline a framework for estimating the aforementioned climate losses using the publicly available ND GAIN dataset, though we emphasize that our approach can be easily adapted to consider other definition of climate losses, such as the ones of Bressan et al. (2024). Furthermore, our model can be extended to include more detailed considerations at corporate level, such as assessing collateral deterioration or other channels of physical risks, particularly relevant for sectors like banking.
@@ 2 The firm value model
$d\left(\log\frac{V_{j}(t)}{V_{j}(0)}\right)=\left(r_{j}-\frac{\sigma^{2}_{j}}{2}-\frac{\omega_{j}^{2}}{2}\right)dt+\sigma_{j}dW_{j}(t)+\omega_{j}dZ(t)$
@@
[…]
@@
We include these two drivers by dividing the firms into clusters with similar climate risk exposures. Each firm $j$in $J$ belongs to a climate cluster $k$in $K$ which contains all companies with similar asset intensity and similar geographical exposures. We model the dynamic of the climate-stressed asset value $\tilde{V}_{j}(t)$ of firm $j$ in cluster $k$ as
@@
where $L_{k}(t)$ is a cluster-specific Compound Poisson process with only negative jumps that represents the climate physical losses on the firm’s assets in the climate-stressed scenario. The Compound Poisson process $L_{k}(t)$ of cluster $k$ is defined as
@@
To evaluate the equity value of company $j$ in the climate stressed scenario we cannot use the Black-Scholes formula as in (4) to price the call option since the firm’s value is a jump-diffusion process. Instead, we utilize the Lewis (2001) formula. The equity value, in the climate-stressed scenario is
@@
[…]
@@
$z_{j}=\log\frac{V_{j}(0)}{D_{j}}+(r_{j}-\gamma_{k})T$ and $\gamma_{k}=\lambda_{k}(1-e^{-\theta_{k}})$
When analyzing portfolio losses in a climate stressed scenario we consider the following simulation framework. We simulate the jump-diffusion process in (5) on a discrete time grid. We simulate the increments of the Gaussian market factor $Z(t)$ and the idiosyncratic factor $W_{j}(t)$. For the jump component at the cluster level, $L_{k}(t)$, we simulate the increments of the Poisson process. The detailed description of the simulation algorithm for the jump-diffusion process (with and without jumps) is available in Algorithm 1. In our numerical analysis, we consider $10^{5}$ simulations and a monthly time-step.
@@ 3 Dataset Composition
To evaluate the physical risk in different geographical areas, we consider the vulnerability score component provided by the Notre-Dame Global Adaptation Initiative (ND-GAIN). [^8] For further detail refer to: https://gain.nd.edu/our-work/country-index/. We utilize the dataset available on December 2023. The vulnerability score reports a country’s exposure, sensitivity, and capacity to adapt to the negative effects of climate change. ND-GAIN aims to quantify the overall vulnerability by considering six life-supporting sectors – food, water, health, ecosystem service, human habitat, and infrastructure.
@@ 4.1 Climate Clustering
4.2 Expected Loss in a climate stressed scenario
@@ 4.3 FVM estimation
Table 4: Comparison between the climate stressed expected loss for each cluster (reported in round brackets) with the one induced by the calibrated model. The values are expressed in absolute value as percentages.
@@
Table 5: Comparison between the available companies in each cluster at the end of climate clustering (reported in round brackets) with the one available for the FVM calibration.

### v2

# Physical Climate Risk in Asset Management

@@ Abstract
Climate-related phenomena are increasingly affecting regions worldwide, manifesting as floods, water scarcity, and heat waves, significantly impairing companies’ assets and productivity. It is essential for asset managers to quantify the exposure of their portfolios to such risk. To this aim, we develop a framework, based on the Vasicek model for credit risk, that introduces downward jumps due to climate phenomena in a company asset’s dynamics. These negative shocks are designed to mirror the negative effect of extreme climate events. The model calibration relies on companies’ asset intensity and geographical exposure. We apply the new multivariate firm value model with jumps to assess the impact of climate-related extreme events on expected and unexpected portfolio losses. Our findings indicate that expected losses increase over time, with pronounced differences in exposure observed across sectoral indices. From an environmental policy perspective, these results suggest the need for additional capital buffers to offset losses arising from physical climate risks, particularly in sectors with high asset intensity.
$(\ddagger)$ Politecnico di Milano, Department of Mathematics, 32 p.zza L. da Vinci, Milano
$(+)$ CREST, ENSAE, Institut Polytechnique de Paris, 5 Avenue Henry Le Chatelier, Palaiseau
()(*) Arca Fondi SGR, 3, Via Disciplini, Milano
@@
Running title: Physical Climate Risk in Asset Management
@@ 1 Introduction
Physical risk emerging from climate change will play a key role in the asset management and banking industry. It comprises both acute risks stemming from climate-related hazards (e.g., from extreme weather events) and chronic risks (e.g., from sea-level rise) from long-run climate impacts (Chou et al. 2023, De Giuli et al. 2024, Bressan et al. 2024). In response, central banks and supervisory authorities globally expect financial institutions with material climate-related and environmental exposures to integrate physical climate risks into their risk management frameworks. [^1] See e.g., the ECB https://www.ecb.europa.eu/press/economic-bulletin/focus/2023/html/ecb.ebbox202302_06~0e721fa2e8.en.html, the FED https://www.federalreserve.gov/publications/files/csa-exercise-summary-20240509.pdf, and the Basel Committee on Banking Supervision perspectives https://www.bis.org/bcbs/publ/d560.pdf. We propose a new and simple model to include this risk in asset management decisions, using a multivariate approach. We model physical risk impact on stock portfolios, aggregating the firm-level effects of such risk. To the best of our knowledge, this is the first model in the literature that exhibits these characteristics and that can be flexibly applied to any stocks portfolio using widely available data.
@@
Physical risk affects directly firms’ tangible assets (e.g. production plants or equipment) as well as operating costs, such as relocating and insurance costs. These shocks result in lower profits and reduced repayment capacity, which are then transmitted to the listed stock prices (Ginglinger and Moreau 2023). To mirror this cascading effect we base our approach on Vasicek (1987), a multivariate generalization of the simple Merton (1974) Firm Value Model (FVM) that considers a single market risk factor. In this framework, the firms’ asset value is modeled with a geometric Brownian motion, while equity is modeled as a European call option on the firm’s asset value, with the company’s debt as strike price. We propose a modification of the Vasicek (1987) model that incorporates jumps to capture the impact of climate events which adversely affect a firm’s asset value. The jump process frequency and size depend on the firm’s characteristics such as its geographic exposure and how much of its revenues are linked to physical assets.
@@
In our study, we follow an efficient and versatile approach for the calibration. It is based on widely available data that covers most listed stocks worldwide and enables low-cost implementation by financial institutions. Specifically, we focus on the exposure of physical assets to climate risks. We utilize the publicly available ND-GAIN (Notre Dame Global Adaptation Initiative) Country Index to assess risk based on geographical location and incorporate balance sheet data to estimate the proportion of a company’s assets exposed to such risks. [^2] See https://gain.nd.edu/assets/522870/nd_gain_countryindextechreport_2023_01.pdf. The ND-GAIN index is a widely used benchmark for climate risk in the literature (see e.g., Russo et al. 2019, Dechezleprêtre et al. 2022). However, our model is very flexible since it can be calibrated starting from different definitions of climate losses such as the ones computed by Bressan et al. (2024) for Mexican companies or from private data providers such as those listed in Hain et al. (2022) or the CRIS methodology discussed in Ginglinger and Moreau (2023).
@@
Once the extended Vasicek (1987) model is calibrated, we can utilize it to compute different climate-related risk metrics for a portfolio. In our analysis, we focus on computing an add-on for the Value-at-Risk (VaR) of a portfolio that accounts for a climate-stressed scenario. We then evaluate the evolution of such metrics for different time horizons. This is particularly relevant from the perspective of financial institutions, as it enables the incorporation of physical risks into their risk management practices building upon their existing frameworks. Moreover, from an environmental policy perspective, these findings suggest that regulators and supervisors should introduce additional capital requirements for exposed portfolios to offset such risks.
Climate Change Implications
The Intergovernmental Panel on Climate Change (IPCC) Synthesis Report confirms that more than a century of fossil fuel combustion, along with unsustainable and unequal energy and land use, has unequivocally caused global warming. Between 2011 and 2020, the global surface temperature reached 1.1°C above 1850–1900 levels, resulting in widespread adverse impacts and losses for both people and ecosystems. The decade from 2015 to 2024 was the warmest on record, with 2024 being the warmest year and 2016 being the second warmest. Six global temperature datasets show the five-year average (2020-2024) as the highest on record. The ten warmest years in the 175-year record have all occurred within this decade. Current nationally determined contributions imply that global warming will reach 1.5°C in the early 2030s, significantly reducing the likelihood of limiting temperature rise to 2.0°C by the end of the century. Each additional increment of warming will intensify multiple and overlapping climate-related hazards across all world regions.
Consequently, the last decade saw a sharp increase in extreme weather events, such as heatwaves, floods, and wildfires, leading to unprecedented economic losses, with several disasters exceeding USD 10 billion. Ice sheet loss in the Arctic and Antarctic accelerated, contributing to rising sea levels. CO2 concentrations reached record highs, further intensifying climate change. In recent years, wildfire activity across the globe has reached unprecedented levels, causing record-breaking burn areas and carbon emissions, particularly in regions such as Canada, Amazonia, and the Mediterranean, where extreme events were estimated to be up to 20 times more likely than in a pre-industrial climate (Jones et al. 2024). The impacts of climate change have disproportionately affected the most vulnerable people and systems, particularly in least developed countries in the Global South, despite their low per capita emissions. Vulnerability is highest in regions with poverty, limited services, weak governance, conflict, and climate-sensitive livelihoods, such as parts of Africa, South Asia and Central and South America. Human and ecosystem vulnerabilities are deeply interconnected and shaped by inequity, unsustainable practices, and historical injustices (Damert and Baumgartner 2018).
Despite improved early warning systems, climate-related displacement surged, highlighting the growing risks and the urgent need for climate adaptation and mitigation strategies. Numerous initiatives have been introduced by various countries and international organizations - such as the United Nations Framework Convention on Climate Change, the United Nations Environment Programme, and the IPCC - to address global climate change through both mitigation and adaptation strategies. A mitigation strategy refers to actions taken to reduce or prevent the emission of greenhouse gases into the atmosphere, including activities such as carbon sequestration, the Clean Development Mechanism, joint implementation, as well as the promotion of renewable and non-polluting energy sources such as solar, wind, and geothermal power (Wang et al. 2023). On the other hand, adaptation strategies include the processes of adjustment to actual or expected climate changes and its effects, in order to moderate harm or exploit beneficial opportunities. [^3] For a comprehensive analysis of climate change mitigation and adaptation strategies, we refer the reader to the IPCC Sixth Assessment Report (2022), available at: https://www.ipcc.ch/report/ar6/wg2/downloads/report/IPCC_AR6_WGII_SummaryVolume.pdf These adverse effects can have severe financial consequences for firms that can be adversely affected if ill-prepared (Gasbarro and Pinkse 2016).
Risk and Climate Risk
Risk, in general, refers to the uncertainty surrounding future outcomes, encompassing both the possibility of loss and gain. In financial contexts, risk specifically denotes the potential for monetary loss due to adverse movements in financial variables such as interest rates, exchange rates, or asset prices (McNeil et al. 2015). Financial risk is multifaceted and includes market risk, credit risk, liquidity risk, and operational risk, each of which can significantly affect the stability and performance of firms and financial systems. Understanding and quantifying these risks is central to financial decision-making and regulatory frameworks. As noted by Jorion and Zhang (2007), effective risk management requires a systematic approach to identifying, measuring, and mitigating exposures in an increasingly complex and interconnected financial environment.
A new class of risk that spans all aforementioned traditional categories of financial risk is climate risk. Climate risk is commonly categorized into two main components: physical risk and transition risk. Physical risk refers to the potential economic and financial losses that arise from the physical impacts of climate change events. These risks stem from both acute events (such as hurricanes, floods, wildfires, and heatwaves) and chronic changes (such as rising sea levels, long-term temperature increases, and shifts in precipitation patterns). The first class represents short-in-time but severe events, while chronic changes are identified by longer-term shifts in climate patterns. We can define climate transition risk as the unanticipated financial gains or losses arising from a faster-than-expected transition toward a carbon-neutral economy. It includes risks created by mitigation and adaptation policies, emerging clean technologies, and behavioral changes of consumers and investors (Bolton and Kacperczyk 2023, Fliegel 2025): some firms will benefit from the shift to a low-carbon economy, while others, particularly those heavily reliant on fossil fuel-based activities, will incur significant losses.
In this paper, we focus on the physical climate market risk, i.e. how extreme weather phenomena could affect a portfolio of stocks. Physical climate risks affect asset values through several key transmission channels. Acute physical events, such as floods, hurricanes, or wildfires, can lead to the destruction of physical assets, disruption of operations and supply chain, and increased insurance costs. Chronic risks, including rising temperatures and sea-level rise, can gradually reduce land usability, agricultural productivity, and labor efficiency. These physical impacts can translate into lower revenues, higher operating costs, or even stranded assets for affected firms. In turn, this deteriorates firms’ financial performance and creditworthiness, leading to downward revisions in equity valuations and increased default risk on debt instruments (see Bressan et al. 2024, and references therein). At the systemic level, widespread physical damages can trigger portfolio reallocation, risk repricing, and higher market volatility, especially in sectors or regions with high exposure to climate hazards.
Asset Management in a Climate Change Scenario
The aim of this work is to develop a reliable model for assessing physical climate market risk within the context of asset management. In particular, we focus on asset managers, broadly defined as financial institutions or professional entities that manage investments on behalf of clients—including individuals, pension funds, insurance companies, and sovereign wealth funds—with the goal of optimizing returns for a given level of risk. Within this framework, we consider financial firms that invest in diversified portfolios of equities. The objective is to quantify how physical risks stemming from climate change - such as extreme weather events, temperature anomalies, or rising sea levels - can propagate through market channels, impacting the value and risk profile of portfolios under management.
@@
The majority of the literature on climate risk impact on portfolios focuses on transition risk and not on physical risk. Le Guenedal and Tankov (2024) utilize a FVM that accounts for the uncertainty of the carbon tax trajectory while Kölbel et al. (2024) show that, in a FVM approach, CDS spread are affected by transition and physical risk news. De Angelis et al. (2023) discuss the influence of green investors on long-run emissions trajectories, finding that tighter climate regulations and climate-related technological innovations abate carbon emissions, in contrast to uncertainty regarding future climate. Bolton and Kacperczyk (2023) point out that transition risk is not significantly related to different exposures to physical risk.
@@
Main contributions
This paper adds to the existing literature by proposing an alternative approach for physical risk in asset management that strikes a balance between the two extrema discussed above: Dietz et al. (2016) estimate only a global loss, while Bressan et al. (2024) provide highly detailed asset-level estimations for a single country, which are challenging to apply for portfolio risk assessment. Our model, in contrast, can be easily applied to any portfolio of assets for which climate loss estimates and balance sheet data are available, providing various climate-related risk metrics at both the portfolio and asset levels. We also outline a framework for estimating the aforementioned climate losses using the publicly available ND-GAIN dataset, though we emphasize that our approach can be easily adapted to consider other definition of climate losses, such as the ones of Bressan et al. (2024). Furthermore, our model can be extended to include more detailed considerations at corporate level, such as assessing collateral deterioration or other channels of physical risks, particularly relevant for sectors like banking.
@@ 2 The firm value model
$\hskip-39.83368ptd\left(\log\frac{V_{j}(t)}{V_{j}(0)}\right)=\left(r_{j}-\frac{\sigma^{2}_{j}}{2}-\frac{\omega_{j}^{2}}{2}\right)dt+\sigma_{j}dW_{j}(t)+\omega_{j}dZ(t)$
@@
$\begin{aligned} \omega_{j}&=:\hat{\sigma}_{j}\sqrt{\rho}\\ \sigma_{j}&=:\hat{\sigma}_{j}\sqrt{1-\rho}\end{aligned}\;\;,$ (3)
@@
We include these two drivers by dividing the firms into clusters with similar climate risk exposures. Each firm $j$ $\in$ $J$ belongs to a climate cluster $k$ $\in$ $K$ which contains all companies with similar asset intensity and similar geographical exposures. We model the dynamic of the climate-stressed asset value $\tilde{V}_{j}(t)$ of firm $j$ in cluster $k$ as
@@
where $L_{k}(t)$ is a cluster-specific Compound Poisson process with only negative jumps that represents the physical climate losses on the firm’s assets in the climate-stressed scenario. The Compound Poisson process $L_{k}(t)$ of cluster $k$ is defined as
@@
To evaluate the equity value of company $j$ in the climate-stressed scenario we cannot use the Black-Scholes formula as in (4) to price the call option since the firm’s value is a jump-diffusion process. Instead, we utilize the Lewis (2001) formula. Lewis’ approach (based on Fourier transform methods) allows pricing in general settings, including stochastic volatility and jump processes, in particular when the characteristic function of the log-process is known. While Black–Scholes yields simple analytical formulas under strong assumptions, Lewis provides a flexible framework to derive option prices semi-analytically from the characteristic function of the underlying process. From a computational point of view, the Lewis formula is comparable to the Black and Scholes one because it involves just an integration in the complex plane. The equity value, in the climate-stressed scenario is
@@
$=V_{j}(0)e^{-\gamma_{k}T}\left[1-\frac{e^{-\frac{z_{j}}{2}}}{2\pi}\int_{-\infty}^{+\infty}\frac{e^{-ivz_{j}}\Phi(-v-\frac{i}{2})}{v^{2}+\frac{1}{4}}dv\right]$ (7)
@@
$z_{j}=\log\frac{V_{j}(0)}{D_{j}}+(r_{j}-\gamma_{k})T$ and $\gamma_{k}=\lambda_{k}(1-e^{-\theta_{k}})$.
When analyzing portfolio losses in a climate-stressed scenario we consider the following simulation framework. We simulate the jump-diffusion process in (5) on a discrete time grid. We simulate the increments of the Gaussian market factor $Z(t)$ and the idiosyncratic factor $W_{j}(t)$. For the jump component at the cluster level, $L_{k}(t)$, we simulate the increments of the Poisson process. The detailed description of the simulation algorithm for the jump-diffusion process (with and without jumps) is available in Algorithm 1. In our numerical analysis, we consider $10^{5}$ simulations and a monthly time-step.
@@ 3 Dataset Composition
To evaluate the physical risk in different geographical areas, we consider the vulnerability score component provided by the Notre-Dame Global Adaptation Initiative (ND-GAIN). [^8] For further detail refer to: https://gain.nd.edu/our-work/country-index/. We utilize the dataset available on December 2023. The vulnerability score reports a country’s exposure, sensitivity, and capacity to adapt to the negative effects of climate change. ND-GAIN aims to quantify the overall vulnerability by considering six life-supporting sectors – food, water, health, ecosystem service, human habitat, and infrastructure. We report the dataset descriptive statistics in Appendix A.
@@ 4.1 Climate Clustering
4.2 Expected Loss in a climate-stressed scenario
@@ 4.3 FVM estimation
Table 4: Comparison between the climate stressed expected loss for each cluster (reported in round brackets) with the one induced by the calibrated model. The values are expressed in absolute value as percentages. The rows correspond to vulnerability clusters, while the columns represent intensity clusters.
@@
Table 5: Comparison between the available companies in each cluster at the end of climate clustering (reported in round brackets) with the one available for the FVM calibration. The rows correspond to vulnerability clusters, while the columns represent intensity clusters.